Hi, ive been asked to answer the following logarithm questions, with great difficulty, there doesnt seem any decent books out there, im just wondering if you could possibly help.
1) 3log(x-2) + 1/2log(x) ( that is half log(x) )
and
2) log(3xy) - log(x)
there is no "already answered questions" and i would be most gratefull of the answer or idea for the answer.
There is also another question which i would like your opinion and solution on,
it is
1) (5+j3)(3-j)-(3+j)
I have had many attempts at this coming up with such answers as
12 + j2 and 15 + j3 , am i on the right lines of thinking or completely wrong.
I am looking forward to you results and opinions which i will highly regard.
Thank you very much for your time.
regards
pete

Answer by stanbon(48535)

1) 3log(x-2) + 1/2log(x) ( that is half log(x) )
Usually the instructions for a problem like this is to
simplify this log expression. If that is the case here
the answer would be:
=log[(x-2)^2*(sqrtx)]
=log [(x^2-4x+4)*(sqrtx)]
--------------
2) log(3xy) - log(x)
=log[3xy/x]
=log(3y)
---------------
There is also another question which i would like your opinion and solution on,
it is
1) (5+j3)(3-j)-(3+j)
"j" usually means sqrt(-1)
If that is so for your problem you would get:
=(15-5j+9j+3)-(3+j)
=(18+4j)-(3+j)
=15+3j
-------
Hope this helps.
Cheers,
Stan H.